OpenStudy (anonymous):

An arc of 1 inch is part of a circle of radius 1 inch. What is the measure of the central angle that forms this arc? 57.3 B. 60 C. 90 D. 14.6

OpenStudy (anonymous):

can u sketch it out?

Parth (parthkohli):

No need to sketch this. Start by finding out the circumference. \[2 \times \pi\times 1 \implies 2\pi\]

Parth (parthkohli):

Also remember that an arc is directly proportionate with the central angle. That can result in a proportion that looks like this:\[2\pi :360 = 1:x \]

OpenStudy (anonymous):

so its 2*3.14 dividedby 360/x

OpenStudy (anonymous):

ahh dnt get it

OpenStudy (anonymous):

57 degree

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

tafara27